# Bayesian Intrinsic Groupwise Image Registration: Unsupervised Disentanglement of Anatomy and Geometry

### v1

@@ 3.1 Hierarchical Bayesian Inference
In practice, registration is often performed at multiple levels to facilitate convergence of the algorithm [71, 72]. Therefore, we express the latent variables with $L$ hierarchical levels, i.e. $\bm{z}=(\bm{z}^{l})_{l=1}^{L}$ and $\bm{v}_{j}=(\bm{v}_{j}^{l})_{l=1}^{L}$, in which higher levels indicate finer resolutions. Thus, the graphical model of the image generative process can be described as Fig. 0(a). Note that while different levels of the latent variables are assumed to be independent, the inference procedure is performed hierarchically.
Since we are going to parameterize the likelihood $p(\bm{u}\mid\bm{z},\bm{v};\bm{\theta})$ by neural networks in the context of Bayesian deep learning, exact inference and parameter estimation becomes intractable [56]. Therefore, we resort to variational inference (VI) to approximate the maximum likelihood. The objective function of VI is the evidence lower bound (ELBO) of the log-likelihood function, i.e.
@@
where $q(\bm{v}\mid\bm{u};{\bm{\psi}})=\prod_{l=1}^{L}q(\bm{v}^{l}\mid\bm{u},\bm{v}^{<l};{\bm{\psi}})$, with $\bm{v}^{l}\triangleq(\bm{v}_{j}^{l})_{j=1}^{N}$, and $\bm{v}^{<l}$ denotes the velocity fields in levels lower than $l$. This factorization over velocity fields can be illustrated by the inference steps #1 in Fig. 0(b). Besides, we assume that the multi-level variational posteriors of the velocity fields factorizes as
@@
=\ = $\sum_{j=1}^{N}\sum_{l=1}^{L}\mathbb{E}_{q(\bm{v}_{j}^{<l}\mid\bm{u};{\bm{\psi}})}\left[D_{\text{KL}}\big{[}q(\bm{v}_{j}^{l}\mid\bm{u},\bm{v}^{<l};{\bm{\psi}})\;\|\;p(\bm{v}_{j}^{l})\big{]}\right].$
@@
=\ = $\sum_{l=1}^{L}\mathbb{E}_{q(\bm{z}^{<l}\mid\bm{u},\bm{v};{\bm{\psi}})}\left[D_{\text{KL}}\big{[}q(\bm{z}^{l}\mid\bm{u},\bm{v};{\bm{\psi}})\;\|\;p(\bm{z}^{l})\big{]}\right],$
@@
Eq. 11 summarizes the decomposition of KL divergence w.r.t. the latent variables. Note that for simplicity we have defined $q(\bm{v}_{j}^{<1}\mid\bm{u};{\bm{\psi}})=q(\bm{z}^{<1}\mid\bm{u},\bm{v};{\bm{\psi}})\triangleq 1$. As will be elucidated in the following subsections, the key idea behind Eq. 11 is: the overall KL divergence is decomposed w.r.t. (i) the common structural representations $\bm{z}$, and (ii) the velocity fields $\bm{v}$. The former is used to measure the intrinsic structural dissimilarity among the observed multi-modal images, while the latter serves as smoothness regularization to enforce diffeomorphism of the transformations.
@@ 3.2 Intrinsic Distance over Structural Representations
As mentioned before, the KL divergence w.r.t. the common structural representations is used to measure the dissimilarity of the observed images for registration. Namely, instead of using similarity measures in the image space, we propose to learn an intrinsic distance over the structural representations corresponding to the observed images.
We first extract the single-view structural representation maps corresponding to each individual image, represented by the distributions $q_{j}^{\diamond}(\bm{z}\mid u_{j},\bm{v}_{j};{\bm{\psi}_{j}})$, $j=1,\dots,N$, which is predicted from the inference steps #2 (ref. Fig. 0(c)). Note that $\bm{\psi}_{j}$ denotes the modality-specific variational parameters for each image, which are the same for images of the same modality. Then we define the variational posterior of the common anatomy as the geometric mean of these single-view posteriors, i.e.
@@
which captures the common information among the multi-modal images. On the other hand, the prior distribution of the common anatomy is given by the arithmetic mean of the single-view posteriors, i.e.
@@
which expresses a priori knowledge of the common anatomy by the mixture of experts. This setup is closely related to the variational mixture of posteriors prior (VampPrior) introduced in [67, 68], where the VampPrior with pseudo-inputs is chosen to maximize the ELBO. For example, Fig. 2 illustrates the geometric and arithmetic mean of Gaussian distributions when $N=3$. One can observe that the three individual modes of single-view posteriors collapse to one after computing the geometric mean. Indeed, if the experts are Gaussian, one can prove that the KL divergence from the arithmetic to the geometric mean distribution is minimized when the experts are identical.
@@
where $\widetilde{D}_{S}$ is an upper bound of the original (possibly) intractable KL divergence $D_{S}$ involving mixture distributions. Thus, in practice we use this upper bound as a surrogate for minimizing the KL divergence w.r.t. the latent variable $\bm{z}$.
Note that when minimizing $\widetilde{D}_{S}$ w.r.t. the distribution $q(\bm{z}^{l}\mid\bm{u},\bm{v};{\bm{\psi}})$, we also obtain the geometric mean $q^{*}(\bm{z}^{l}\mid\bm{u},\bm{v};{\bm{\psi}})$ in Eq. 12, i.e.
@@
where $\widetilde{D}_{S}[q]$ is regarded as a functional of the distribution $q$. This also justifies the usage of geometric mean as the variational posterior of the common anatomy. In fact, as KL divergence is a statistical distance between probability distributions, the $\widetilde{D}_{S}$ in Eq. 14 serves as an intrinsic distance between the structural representations of the common anatomy and each observed image. Therefore, since geometric mean is the distribution that minimizes this intrinsic distance, the optimization of $\widetilde{D}_{S}^{*}\triangleq\widetilde{D}_{S}[q^{*}]$ w.r.t. $q_{j}^{\diamond}$’s will then seek to force the single-view structural representations $q_{j}^{\diamond}(\bm{z}^{l}\mid u_{j},\bm{v}_{j};{\bm{\psi}_{j}})$’s (which are conditioned on the transformation variables) to be identical, thus driving the registration process. Appendix A provides more details for the above reasoning.
@@ 3.2.1 Structural Representations for Variable Group Sizes
In our previous work [73], we have used Gaussian distributions to parameterize the structural representations. That is, we assume $q_{j}^{\diamond}(\bm{z}\mid u_{j},\bm{v}_{j};{\bm{\psi}_{j}})=\mathcal{N}(\bm{z};\bm{\mu}_{\bm{z},j}^{\diamond},\bm{\Sigma}_{\bm{z},j}^{\diamond})$ where $\bm{\Sigma}_{\bm{z},j}^{\diamond}$ are diagonal, and for notational conciseness we omit the superscript $l$ indicating the level of the latent variables. We can calculate that the geometric mean is also a Gaussian distribution, i.e. $q^{*}(\bm{z}\mid\bm{u},\bm{v};{\bm{\psi}})=\mathcal{N}(\bm{z};\bm{\mu}_{\bm{z}}^{*},\bm{\Sigma}_{\bm{z}}^{*})$, where
@@
Therefore, the intrinsic distance has a closed-form expression, i.e.
@@
=\ = $\frac{1}{2}\Bigg{[}\log\frac{\absolutevalue{\bm{\Sigma}_{\bm{z},j}^{\diamond}}}{\absolutevalue{\bm{\Sigma}_{\bm{z}}^{*}}}+\operatorname{tr}\left(\bm{\Sigma}_{\bm{z},j}^{\diamond\,-1}\bm{\Sigma}_{\bm{z}}^{*}\right)$
@@
Here, we extend the framework using categorical distribution to improve interpretability of the structural representations. Particularly, we propose to model them by independent categorical latent variables, i.e., $q_{j}^{\diamond}(\bm{z}\mid u_{j},\bm{v}_{j};{\bm{\psi}_{j}})=\operatorname{Cat}(\bm{z};\bm{\pi}_{j,1}^{\diamond},\dots,\bm{\pi}_{j,K}^{\diamond})$, where $\bm{\pi}_{j,k}^{\diamond}=(\pi_{\bm{\omega},j,k}^{\diamond})_{\bm{\omega}\in\Omega}\in[0,1]^{\absolutevalue{\Omega}}$ and $\sum_{k=1}^{K}\bm{\pi}_{j,k}^{\diamond}=\bm{1}\in\mathds{R}^{\absolutevalue{\Omega}}$. Thus, the geometric mean becomes another categorical distribution
@@ 3.2.2 Categorical Reparameterization Using Gumbel-Rao
Specifically, for objective functions like the ELBO, the gradient w.r.t. the distribution parameters of an expectation over a function $f(\bm{z})$ needs to be computed, namely $\nabla_{\bm{\psi}}\mathbb{E}_{q(\bm{z};{\bm{\psi}})}[f(\bm{z})]$, where $f(\bm{z})$ can be the likelihood function of the observed variables. To overcome the challenge of gradient computation with discrete stochasticity, the Straight-Through Gumbel-Softmax (ST-GS) estimator [75, 77] uses a continuous relaxation, giving rise to a biased Monte-Carlo gradient estimator of the form
@@
where $\bm{G}^{s}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\bm{g}+\log\bm{\pi}\mid\bm{z}$ for $s=1,\dots,S$, and $\bm{g}$ is a random vector with its entries i.i.d. sampled from the Gumbel distribution. More details on the Gumbel-Rao estimator can be found in Appendix B.
@@ 3.3 Spatial Regularization for Diffeomorphisms
To impose spatial smoothness of the velocity fields, we define its prior distribution by $p(\bm{v}_{j}^{l})=\mathcal{N}(\bm{v};\bm{0},\bm{\Lambda_{v}}^{-1})$, where the precision matrix $\bm{\Lambda_{v}}$ is given by the scaled Laplacian matrix of a neighbourhood graph on the voxel grid [46], i.e. $\bm{\Lambda_{v}}=\lambda\bm{L_{v}}=\lambda(\bm{D_{v}}-\bm{A_{v}})$, with $\bm{D_{v}}$ and $\bm{A_{v}}$ the degree and adjacency matrices, respectively. On the other hand, the variational posterior of the velocity fields is given by a mean-field Gaussian distribution, i.e. $q(\bm{v}_{j}^{l}\mid\bm{u},\bm{v}^{<l};{\bm{\psi}})=\mathcal{N}(\bm{v};\bm{\mu}_{\bm{v},j}^{l},\bm{\Sigma}_{\bm{v},j}^{l})$ where $\bm{\Sigma}_{\bm{v},j}^{l}$ is diagonal. Therefore, the KL divergence for the velocity fields has an explicit formula
@@
=\ = $\frac{1}{2}\left[\operatorname{tr}(\lambda\bm{D_{v}}\bm{\Sigma}_{\bm{v},j}^{l}-\log\bm{\Sigma}_{\bm{v},j}^{l})+\frac{\lambda}{2}\sum_{r}\sum_{q\in\mathcal{N}(r)}(\bm{\mu}_{\bm{v},j}^{l}[r]-\bm{\mu}_{\bm{v},j}^{l}[q])^{2}\right]$
@@
4 An Interpretable Registration Model via Bayesian Disentanglement Learning
Figure 3: An overview of the proposed disentangled auto-encoding framework for interpretable groupwise registration.
Figure 4: The network architecture for the proposed Bayesian groupwise registration, composed of the encoders that extract categorical structural representation maps, the registration modules that calculate multi-scale velocity fields, and the decoders that reconstruct the original images based on the common structural representations. Without loss of generality, the illustration is with $L=3$ levels and $N=3$ images to co-register. Note that the inference steps #2 are only performed in the training stage, while in the test stage the encoder is just fed with the original image group to predict groupwise registration. The purple boxes indicate the calculation of related terms in the ELBO.
Most learning-based registration methods in recent literature are based on an end-to-end training pipeline, where a black-box neural network directly predicts the spatial transformation by optimizing image-level similarity measures [44, 45, 46, 78, 79]. However, not only do the end-to-end black-box networks lack interpretability, learning the complex mapping from multi-modal images to their spatial correspondence disregards their underlying structural relationship and may be prone to generalization error [80].

### v2

@@ 3.1 Hierarchical Bayesian Inference
In practice, registration is often performed at multiple levels to facilitate convergence of the algorithm [72, 73]. Therefore, we express the latent variables with $L$ hierarchical levels, i.e., $\bm{z}=(\bm{z}^{l})_{l=1}^{L}$ and $\bm{v}_{j}=(\bm{v}_{j}^{l})_{l=1}^{L}$, in which higher levels indicate finer resolutions. Thus, the graphical model of the image generative process can be described as Fig. 1(a). Note that while different levels of the latent variables are assumed to be independent, the inference procedure is performed hierarchically.
Since we are going to parameterize the likelihood $p(\bm{u}\mid\bm{z},\bm{v};\bm{\theta})$ by neural networks in the context of Bayesian deep learning, exact inference and parameter estimation become intractable [58]. Therefore, we resort to variational inference (VI) to approximate the maximum likelihood. The objective function of VI is the evidence lower bound (ELBO) of the log-likelihood function, i.e.,
@@
where $q(\bm{v}\mid\bm{u};{\bm{\psi}})=\prod_{l=1}^{L}q(\bm{v}^{l}\mid\bm{u},\bm{v}^{<l};{\bm{\psi}})$, with $\bm{v}^{l}\triangleq(\bm{v}_{j}^{l})_{j=1}^{N}$, and $\bm{v}^{<l}$ denotes the velocity fields in levels lower than $l$. This factorisation of velocity fields can be illustrated by the inference steps #1 in Fig. 1(b). Besides, we assume that the multi-level variational posteriors of the velocity fields factorizes as
@@
$=$ $\sum_{j=1}^{N}\sum_{l=1}^{L}\mathbb{E}_{q(\bm{v}_{j}^{<l}\mid\bm{u};{\bm{\psi}})}\left[D_{\text{KL}}\big{[}q(\bm{v}_{j}^{l}\mid\bm{u},\bm{v}^{<l};{\bm{\psi}})\;\|\;p(\bm{v}_{j}^{l})\big{]}\right].$
@@
$=$ $\sum_{l=1}^{L}\mathbb{E}_{q(\bm{z}^{<l}\mid\bm{u},\bm{v};{\bm{\psi}})}\left[D_{\text{KL}}\big{[}q(\bm{z}^{l}\mid\bm{u},\bm{v};{\bm{\psi}})\;\|\;p(\bm{z}^{l})\big{]}\right],$
@@
Eq. 11 summarises the decomposition of KL divergence w.r.t. the latent variables. Note that for simplicity, we have defined $q(\bm{v}_{j}^{<1}\mid\bm{u};{\bm{\psi}})=q(\bm{z}^{<1}\mid\bm{u},\bm{v};{\bm{\psi}})\triangleq 1$. As will be elucidated in the following subsections, the key idea behind Eq. 11 is: the overall KL divergence is decomposed w.r.t. (i) the common structural representations $\bm{z}$, and (ii) the velocity fields $\bm{v}$. The former is used to measure the intrinsic structural dissimilarity among the observed multi-modal images, while the latter serves as smoothness regularisation to enforce diffeomorphism of the transformations.
@@ 3.2 Intrinsic Distance over Structural Representations
As mentioned above, the KL divergence w.r.t. the common structural representations is used to measure the dissimilarity of the observed images for registration. Namely, instead of using similarity measures in the image space, we propose learning an intrinsic distance over the structural representations corresponding to the observed images.
We first extract the single-view structural representation maps corresponding to each individual image, represented by the distributions $q_{j}^{\diamond}(\bm{z}\mid u_{j},\bm{v}_{j};{\bm{\psi}_{j}})$, $j=1,\dots,N$, which is predicted from the inference steps #2 (ref. Fig. 1(c)). Note that $\bm{\psi}_{j}$ denotes the modality-specific variational parameters for each image, which are the same for images of the same modality. Then we define the variational posterior of the common anatomy as the geometric mean of these single-view posteriors, i.e.,
@@
which captures the common information among the multi-modal images. On the other hand, the prior distribution of the common anatomy is given by the arithmetic mean of the single-view posteriors, i.e.,
@@
which expresses a priori knowledge of the common anatomy by the mixture of experts. This setup is closely related to the variational mixture of posteriors prior (VampPrior) introduced in [68, 69], where the VampPrior with pseudo-inputs is chosen to maximise the ELBO. For example, Fig. 2 illustrates the geometric and arithmetic mean of Gaussian distributions when $N=3$. One can observe that the three individual modes of single-view posteriors collapse to one after computing the geometric mean. Indeed, if the experts are Gaussian, one can prove that the KL divergence from the arithmetic to the geometric mean distribution is minimised when the experts are identical.
@@
where $\widetilde{D}_{S}$ is an upper bound of the original (possibly) intractable KL divergence $D_{S}$ involving mixture distributions. Thus, in practice we use this upper bound as a surrogate to minimise the KL divergence w.r.t. the latent variable $\bm{z}$.
Note that when minimising $\widetilde{D}_{S}$ w.r.t. the distribution $q(\bm{z}^{l}\mid\bm{u},\bm{v};{\bm{\psi}})$, we also obtain the geometric mean $q^{*}(\bm{z}^{l}\mid\bm{u},\bm{v};{\bm{\psi}})$ in Eq. 12, i.e.,
@@
where $\widetilde{D}_{S}[q]$ is regarded a functional of the distribution $q$. This also justifies the usage of the geometric mean as the variational posterior of the common anatomy. In fact, since the KL divergence is a statistical distance between probability distributions, the $\widetilde{D}_{S}$ in Eq. 14 serves as an intrinsic distance between the structural representations of the common anatomy and each observed image. Therefore, since the geometric mean is the distribution that minimises this intrinsic distance, the optimisation of $\widetilde{D}_{S}^{*}\triangleq\widetilde{D}_{S}[q^{*}]$ w.r.t. $q_{j}^{\diamond}$’s will then seek to force the single-view structural representations $q_{j}^{\diamond}(\bm{z}^{l}\mid u_{j},\bm{v}_{j};{\bm{\psi}_{j}})$’s (which are conditioned on the transformation variables) to be identical, thus driving the registration process. Section A of the Supplementary Material provides more details for the above reasoning.
@@ 3.2.1 Structural Representations for Variable Group Sizes
In our previous work [74], we have used Gaussian distributions to parameterize the structural representations. That is, we assume $q_{j}^{\diamond}(\bm{z}\mid u_{j},\bm{v}_{j};{\bm{\psi}_{j}})=\mathcal{N}(\bm{z};\bm{\mu}_{\bm{z},j}^{\diamond},\bm{\Sigma}_{\bm{z},j}^{\diamond})$ where $\bm{\Sigma}_{\bm{z},j}^{\diamond}$ are diagonal, and for notational conciseness we omit the superscript $l$ indicating the level of the latent variables. We can calculate that the geometric mean is also a Gaussian distribution, i.e., $q^{*}(\bm{z}\mid\bm{u},\bm{v};{\bm{\psi}})=\mathcal{N}(\bm{z};\bm{\mu}_{\bm{z}}^{*},\bm{\Sigma}_{\bm{z}}^{*})$, where
@@
Therefore, the intrinsic distance has a closed-form expression, i.e.,
@@
$=$ $\frac{1}{2}\Bigg{[}\log\frac{\absolutevalue{\bm{\Sigma}_{\bm{z},j}^{\diamond}}}{\absolutevalue{\bm{\Sigma}_{\bm{z}}^{*}}}+\operatorname{tr}\left(\bm{\Sigma}_{\bm{z},j}^{\diamond\,-1}\bm{\Sigma}_{\bm{z}}^{*}\right)$
@@
Here, we extend the framework using categorical distribution to improve the interpretability of the structural representations. Particularly, we propose to model them by independent categorical latent variables, i.e., $q_{j}^{\diamond}(\bm{z}\mid u_{j},\bm{v}_{j};{\bm{\psi}_{j}})=\operatorname{Cat}(\bm{z};\bm{\pi}_{j,1}^{\diamond},\dots,\bm{\pi}_{j,K}^{\diamond})$, where $\bm{\pi}_{j,k}^{\diamond}=(\pi_{\bm{\omega},j,k}^{\diamond})_{\bm{\omega}\in\Omega}\in[0,1]^{\absolutevalue{\Omega}}$ and $\sum_{k=1}^{K}\bm{\pi}_{j,k}^{\diamond}=\bm{1}\in\mathds{R}^{\absolutevalue{\Omega}}$. Thus, the geometric mean becomes another categorical distribution
@@ 3.2.2 Categorical Reparameterization Using Gumbel-Rao
Specifically, for objective functions like ELBO, the gradient w.r.t. the distribution parameters of an expectation over a function $f(\bm{z})$ must be computed, namely $\nabla_{\bm{\psi}}\mathbb{E}_{q(\bm{z};{\bm{\psi}})}[f(\bm{z})]$, where $f(\bm{z})$ can be the likelihood function of the observed variables. To overcome the challenge of gradient computation with discrete stochasticity, the Straight-Through Gumbel-Softmax (ST-GS) estimator [76, 78] uses a continuous relaxation, giving rise to a biased Monte-Carlo gradient estimator of the form
@@
where $\bm{G}^{s}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\bm{g}+\log\bm{\pi}\mid\bm{z}$ for $s=1,\dots,S$, and $\bm{g}$ is a random vector with its entries i.i.d. sampled from the Gumbel distribution. More details on the Gumbel-Rao estimator can be found in Section B of the Supplementary Material.
@@ 3.3 Spatial Regularization for Diffeomorphisms
To impose spatial smoothness of the velocity fields, we define its prior distribution by $p(\bm{v}_{j}^{l})=\mathcal{N}(\bm{v};\bm{0},\bm{\Lambda_{v}}^{-1})$, where the precision matrix $\bm{\Lambda_{v}}$ is given by the scaled Laplacian matrix of a neighbourhood graph on the voxel grid [46], i.e., $\bm{\Lambda_{v}}=\lambda\bm{L_{v}}=\lambda(\bm{D_{v}}-\bm{A_{v}})$, with $\bm{D_{v}}$ and $\bm{A_{v}}$ the degree and adjacency matrices, respectively. On the other hand, the variational posterior of the velocity fields is given by a mean-field Gaussian distribution, i.e., $q(\bm{v}_{j}^{l}\mid\bm{u},\bm{v}^{<l};{\bm{\psi}})=\mathcal{N}(\bm{v};\bm{\mu}_{\bm{v},j}^{l},\bm{\Sigma}_{\bm{v},j}^{l})$ where $\bm{\Sigma}_{\bm{v},j}^{l}$ is diagonal. Therefore, the KL divergence for the velocity fields has an explicit formula
@@
$=$ $\frac{1}{2}\left[\operatorname{tr}(\lambda\bm{D_{v}}\bm{\Sigma}_{\bm{v},j}^{l}-\log\bm{\Sigma}_{\bm{v},j}^{l})+\frac{\lambda}{2}\sum_{r}\sum_{q\in\mathcal{N}(r)}(\bm{\mu}_{\bm{v},j}^{l}[r]-\bm{\mu}_{\bm{v},j}^{l}[q])^{2}\right]$
@@
4 An Interpretable Registration Architecture via Bayesian Disentanglement Learning
Figure 3: An overview of the proposed interpretable groupwise registration architecture via Bayesian disentanglement learning.
Figure 4: The network architecture for the proposed Bayesian groupwise registration, composed of the encoders that extract categorical structural representation maps, the registration modules that calculate multi-scale velocity fields, and the decoders that reconstruct the original images based on the common structural representations. Without loss of generality, the illustration is with $L=3$ levels and $N=3$ images to co-register. Note that inference steps #2 are performed only in the training stage, while in the test stage the encoder is only fed with the original image group to predict groupwise registration. The purple boxes indicate the calculation of related terms in the ELBO.
Most learning-based registration methods in the recent literature are based on an end-to-end training pipeline, where a black-box neural network directly predicts the spatial transformation by optimising some image-level similarity measures [44, 45, 46, 79, 80]. However, not only do black-box end-to-end networks lack interpretability, learning directly the complex mapping from multi-modal images to their spatial correspondence disregards the underlying structural relationship and may be prone to generalisation issues [81].